OpenStudy (phebe):

What are the solutions to the equation ? 0=x^2+3x-10 A-x=5,2 B-x=-5,-2 C-x=-5,2 D-x=5,-2

OpenStudy (phebe):

@mjoxprm360

OpenStudy (anonymous):

0=x^2+3x-10 (x - 2)(x + 5) = 0 x - 2 = 0, x + 5 = 0 x = 2. x = - 5 x = 2, - 5 answer//

OpenStudy (phebe):

thanx so its C

OpenStudy (anonymous):

mhm

OpenStudy (phebe):

u noit sure

OpenStudy (anonymous):

im sure

OpenStudy (phebe):

ok thanx

OpenStudy (anonymous):

no problem :)